The position of a stone dropped from a cliff is described approximately by x = 5t^2 , where x is in meters and t is in seconds. The +x direction is downwards and the origin is at the top of the cliff. Find the velocity of the stone during its fall as a function of time t.

1
Expert's answer
2022-02-28T13:45:23-0500

We know that gives

"x=5t^2"

Now velocity

"v=\frac{dx}{dt}"

Put value

"v=\frac{d(5t^2)}{dt}=10tm/sec"